# Graph Agent Network: Empowering Nodes with Decentralized Communications Capabilities for Adversarial Resilience

### v1

@@ Abstract
Index Terms:
Graph neural networks, adversarial attack, agent-based model, adversarial resilience.
I Introduction
Graph neural networks (GNNs) have become state-of-the-art models for node classification tasks by leveraging end-to-end global training paradigms to effectively learn and extract valuable information from graph-structured data [1, 2, 3]. However, this approach has also inadvertently introduced inherent vulnerabilities, making GNNs vulnerable to adversarial edge-perturbing attacks. These vulnerabilities arise from GNNs exposing a global end-to-end training interface, which while allowing for precise classification, also provides adversaries with opportunities to attack GNNs. These attacks poses a significant challenge to the practical deployment of GNNs in real-world applications where security and robustness are of paramount importance, leading to a critical issue in various application areas [4, 5, 6, 7, 8], including those where adversarial perturbations undermine public trust [9], interfere with human decision making [10], human health and livelihoods [11].
@@ I Introduction
In response to the edge-perturbing attacks, existing defense mechanisms primarily rely on global-optimization-based defense methods [12]. Their objective is to enhance the robustness of GNNs through adversarial training [13], aiming to increase the tolerant ability of perturbations to defend against potential adversarial perturbations. These approaches may still inherit some inherent vulnerabilities of GNNs, as their model frameworks still follow the global-optimization pattern of GNNs. Representative examples include: (1) RGCN [14] replaces the hidden representations of nodes in each graph convolutional network (GCN) [1] layer to the Gaussian distributions, to further absorb the effects of adversarial changes. (2) GCN-SVD [15] combines a singular value decomposition (SVD) filter prior to GCN to eliminate adversarial edges in the training set. (3) STABLE [16] reforms the forward propagation of GCN by adding functions that randomly recover the roughly removed edges. Unfortunately, the computational universality of GNNs has been recently demonstrated [17, 18], signifying that attributed graphs can be classified into any given label space, even those subject to malicious manipulation. This implies that defense approaches focused on enhancing GNNs’ global robustness [14, 15, 16, 19, 20, 21] remain theoretically vulnerable, with potential weaknesses to targeted secondary attacks. In fact, the vulnerability of existing global-optimization-based defenses has been theoretically proven [22]. Adversaries can reduce the classification accuracy of these defenses once again by launching secondary attacks using these defenses as surrogate models. Experimental evidence demonstrates that even under the protection of state-of-the-art defensive measures [23], secondary attacks targeting these defenses successfully mislead 72.8% of the nodes into making incorrect classifications once again.
@@
To overcome the issues mentioned above, inspired by the natural filtering ability of decentralized intelligence [24, 25, 26, 27], we propose a decentralized agent network called Graph Agent Networks (GAgN) whose principle is shown in Fig. 1. GAgN empowers nodes with autonomous awareness, while limiting their perspectives to their 1-hop neighbors. As a result, nodes no longer rely solely on global-level end-to-end training data. Instead, they progressively gain the perception of the entire network through communication with their neighbors and accomplish self-classification. GAgN is designed to enhance the agent-based model (ABM) [28, 29, 30], ensuring improved compatibility with graph-based scenarios. In GAgN, agents are interconnected within the graph topology and engage in cautious communication to mitigate the impact of potential adversarial edges. As time progresses, agents systematically exchange information to broaden their receptive fields, perceive the global information, and attain decentralized intelligence to thus achieve adversarial resilience.
Specifically, GAgN initialize a graph as an agent network where each agent is endowed with three principal abilities:
  • Storage, which encompasses states representing both the agent’s and 1-hop neighbors’ current features, actions that serve as parameters for inference functions, and the agent’s own label.
  • Inference, which enables an agent to perform global reasoning tasks based on its stored inference functions. These tasks include computing its own embedding, estimating the potential degree of a given node, and determining the neighboring confidence between two specific nodes. The first function equips GAgN with the ability to perform node classification, while the latter two jointly contribute to filtering adversarial edges.
  • Communicate, which enables agents to receive and integrate the states and actions of their 1-hop neighbors to augment their own inferential abilities.
Following this, GAgN instigates communication between agents, initially premised on zero trust. During a communication round, each agent acquires training data received from its neighboring agents through predefined I/O interfaces. These data are then organized into a mini-dataset under a limited view, which is employed to train the agent’s inference function, thereby accomplishing agent-level inference tasks. To generalize this inferential capacity to arbitrary node in the graph, each agent further assimilates inference function parameters from neighboring agents into its own inference function. This integration allows agents for global inference capabilities at the graph-level, even under restricted views.
Finally, after a sufficiently prolonged period of communication (i.e., upon the convergence of the inference function’s loss), GAgN opens new communication channels for agents that adhere to the I/O interface specifications. This prompts mutual inspection among agents to identify and filter out suspect adversarial edges using their well-trained functions.
@@
We demonstrate that GAgN and GraphSAGE [2], currently the most potent GNN, are equivalent in terms of their message passing modes. This finding underpins the node classification capability of GAgN. Concurrently, we rigorously prove that all functions in GAgN can be accomplished by the single-hidden-layer multilayer perceptron (MLPs) [31]. In other words, theoretically, within an agent, only three trainable matrices are required in the trainable part, excluding deep-layer networks, bias matrices, and other trainable parameters, to successfully perform the corresponding tasks. This conclusion offers theoretical backing for substantially reducing GAgN’s computational complexity. In practice, by instantiating nodes as lightweight agents, a GAgN can be constructed to carry out node classification tasks with adversarial resilience.
Our main contributions include the following three aspects:
@@
  • We experimentally show that functions in GAgN are effectively executed, which collectively yielding state-of-the-art classification accuracy in perturbed datasets.
We format the paper as follows: Section II introduces related work. Section III presents the specific technical details of our GAgN. Section IV demonstrates the theoretical results pertinent to GAgN. Section V gives the relevant deployments and results of the experiments. Section VI concludes our work.
Figure 2: The internal structure and communication paradigm of an agent within GAgN. Initially, the agent organizes a miniature dataset based on received data to train its functions. It then integrates the functions of its neighbors to further expand its receptive field of the global graph. Ultimately, the agent trains functions that possess effective inferential capabilities and self-classification.
II Related works
II-A Edge-perturbing Attacks
Edge-perturbing attacks serve as a primary example of adversarial attacks against GNNs. Through an end-to-end training process, the GNN’s message passing mechanism can be strategically altered to manipulate the classification of the target node or diminish the model’s overall performance. Pioneering works in the realm of adversarial attacks on graph structures were proposed by Dai et al.[32] and Zügner et al.[33], leading to the advent of numerous graph attack methodologies specifically tailored for node prediction tasks. Subsequently, in 2019, Aleksandar et al. [34] conducted the inaugural adversarial vulnerability analysis on the broadly employed methods based on random walks and introduced efficient adversarial perturbations that disrupt the network structure. The following year, Chang et al.[35] launched attacks on diverse graph embedding models using a black-box driven method. Concurrently, Wang et al.[36] designed a threat model to characterize the attack surface of collective classification methods, with a focus on adversarial collective classification. The foundation of these attacks resides in the perturbed graph structure, which is the central subject of this article. Ma et al.[37] explored black-box attacks on GNNs under a novel and realistic constraint where attackers can only access a subset of nodes in the network. Moreover, Xi et al.[38] introduced the first backdoor attack on GNNs, significantly elevating the misclassification rate of popular GNNs on real-world data.
II-B Defenses for Edge-perturbing Attacks
On the defensive front, most protection strategies are typically based on the global-optimization patterns. Zügner et al.[39] were the first to propose a method for certifying robustness (or the lack thereof) of graph convolutional networks with respect to perturbations of the node attributes. Zhu et al.[14] put forward RGCNs designed to automatically neutralize the effects of adversarial changes in the variances of Gaussian distributions. Several defense approaches leverage the concept of generation to bolster the model’s robustness. For instance, Deng et al. [40] presented batch virtual adversarial training (BVAT), a novel regularization method for GCNs that enhances model robustness by incorporating perturbed embeddings. Feng [41] proposed Graph Adversarial Training (GAD), which accounts for the influence from connected examples while learning to construct and resist perturbations. They also introduced an adversarial attack on the standard classification to the graph. Tang et al. [42] researched a novel problem of boosting the robustness of GNNs against poisoning attacks by exploiting a clean graph and generating supervised knowledge to train the model in detecting adversarial edges, thus improving the robustness of GNNs. Liao et al.[43] developed a framework to locally filter out pre-determined sensitive attributes through adversarial training with total variation and the Wasserstein distance, providing robust defense across various graph structures and tasks. Liu et al.[19] proposed Elastic GNNs (EGNN) that can adaptively learn representations of nodes and edges on a graph, dynamically adding or removing nodes and edges to defend against malicious edge perturbations. Jin et al.[21] put forth Pro-GNN, which enhances GNN robustness by optimizing a structural loss function. This measure increases the network’s stability and resilience against attacks targeting nodes and edges.
III The Proposed Method
III-A Preliminaries
@@
We consider connected graphs $\mathcal{G}=(\mathcal{V},\mathcal{E})$ consisting $|\mathcal{V}|$ nodes. The degree of node $v_{i}$ is denoted by $\mathrm{deg}_{i}$. The feature vector and on-hot encoding label of node $v_{i}$ is $\mathbf{z}_{i}\in\mathbb{R}^{1\times d_{z}}$ and $\mathbf{\ell}_{i}\in\mathbb{R}^{1\times d_{L}}$, respectively. The maximum degree over nodes in $\mathcal{G}$ is denoted as $\mathrm{deg}_{\mathrm{max}}$. The edge connected $v_{i}$ and $v_{j}$ is written as $e_{i,j}$ or $e_{j,i}$. The neighborhood $\mathrm{N}_{i}$ of a node $v_{i}$ consists of all nodes $v_{j}$ for which $e_{i,j}\in\mathcal{E}$.
@@
III-B Overview
In GAgN, structures of I/O interfaces, computational models, and storage are same for all agents. Despite these similarities, each agent operates independently, given the distinct information it receives from neighboring agents.
The agent comprises the following components:
  • State space that consists of 1) the agent’s and its neighbors’ features, which are communicated through an information exchange mechanism between agents, and 2) inferring results of the degree and the neighbor relationship by given features. Inferring results are typically not computed unless a request for computation is initiated.
  • Action spaces that defines the inferring functions, which including individual and communicable functions.
  • I/O interface that serves as a conduit for data exchange between agents and their neighbors. It defines the types of information that can be exchanged including states and communicable functions.
  • Decision rules that retrieve the parameters of action spaces based on the agent’s current state.
  • Action update rules that aggregate neighboring agents’ actions to update the agent’s itself actions.
Figure 2 illustrates the internal structure and communication paradigm of an agent within GAgN. The state of each agent primarily consists of features utilized for inference and the corresponding inference results, with their format predominantly determined by other inference components. Thus, we initiate our discussion with action spaces, and intersperse the introduction of the state’s structure and update strategies while elaborating on other components, aiming to clearly elucidate the entire process.
III-C Action Spaces
The action space comprises the set of all potential actions that a node can undertake in a given reasoning task [44]. In GAgN, the action space is devised as a series of perception functions, sharing identical parameter structures but different specific parameters. The optimal function parameters of agents for each communication round can be found within the action space. Below, we use node $v_{i}$ as an example to expound on these functions it encompasses.
III-C1 Individual Function
Parameters of individual function are agent-specific, reflecting the understanding they gain through interactions with their neighbors, without necessitating direct agent interactions. This includes:
[…]
$\mathcal{A}_{i}(\mathrm{N}_{i})=\frac{\mathbf{w}_{i}\mathrm{Concat}_{r}(\{\mathbf{z}_{i}\}\cup\{\mathbf{z}_{j}:v_{j}\in\mathrm{N}_{i}\})}{\mathrm{deg}_{i}+1}.$ (1)
Here $\mathbf{w}_{i}$ is initialized to a vector with the smallest possible values, and it is updated iteratively when node $v_{i}$ receives information.
III-C2 Communicable Functions
@@
Notably, while intuitive perception suggests that $\mathcal{D}_{i}$ and $\mathcal{N}_{i}$ can be effectively trained end-to-end, this is infeasible in GAgN due to potential adversaries exist and agents’ limited view, restricting access to ample global training samples.
III-D I/O Interface
The internal structure of a node supports its input and output functionality.
III-D1 Input
In a single communication round, node $v_{i}$ receives the following messages from its neighbors $\mathrm{N}_{i}$. 1) Communicable functions of all nodes in $\mathrm{N}_{i}$, i.e., sets of functions $\mathcal{M}_{i}^{\mathrm{rec}}=\{\mathcal{M}_{j}:v_{j}\in\mathrm{N}_{i}\}$, $\mathcal{D}_{i}^{\mathrm{rec}}=\{\mathcal{D}_{j}:v_{j}\in\mathrm{N}_{i}\}$, and $\mathcal{N}_{i}^{\mathrm{rec}}=\{\mathcal{N}_{j}:v_{j}\in\mathrm{N}_{i}\}$. 2) Features from the 1-hop neighbors $Z^{\mathrm{I}}_{i}=\{\mathbf{z}_{j}:v_{j}\in\mathrm{N}_{i}\}$. 3) Sampled features from the 2-hop neighbors $Z^{\mathrm{II}}_{i}=\bigcup_{v_{j}\in\mathrm{N}_{i}}\mathrm{S}(v_{j})-Z^{\mathrm{I}}_{i}$, where $\mathrm{S}(\cdot)$ denotes the sample function. Note that second-order neighbor features are transmitted through immediate neighbors, so the sampling $\mathrm{S}(\cdot)$ occurs when the corresponding first-order neighbor sends messages out.
III-D2 Output
@@
where $\mathrm{rank}(\cdot)$ denotes the rank, i.e., a smaller rank indicating a larger value. 3) Communicable functions $\mathcal{M}_{i}$, $\mathcal{D}_{i}$, and $\mathcal{N}_{i}$.
III-E Decision Rules
@@
The decision rules dictate the mechanism for choosing the corresponding action in different states [45, 46]. In GAgN, the decision rules are specified as the mechanism for training parameters of the functions in the action space, which is based on $v_{i}$’s states after receiving messages from its neighbors. These functions compute the adjustment direction for their parameters through the loss function and obtain trained parameters using stochastic gradient descent (SGD). Next, we introduce these loss-based rules separately.
III-E1 Decision Rule of $\mathcal{A}_{i}$
The loss function $J_{\mathcal{A}}(\cdot)$ of $\mathcal{A}_{i}$ encourages to assign different weights to the neighbors of $v_{i}$, with the aim of obtaining a more accurate embedding that is closer to its label. Therefore, given a aggregated feature $\mathbf{z}^{\mathcal{A}}_{i}=\mathcal{A}_{i}(\mathrm{N}_{i})$, based on the cross-entropy loss,
$J_{\mathcal{A}}(z^{\mathcal{A}}_{i})=-\log\mathcal{M}_{i}(\mathbf{z}^{\mathcal{A}}_{i})\mathbf{\ell}_{i}^{\top}.$ (3)
Then, the parameters of $\mathcal{A}_{i}$ can be updated through stochastic gradient descent based backpropagation. It should be noted that although $\mathcal{M}_{i}$ is involved in the computation during the forward propagation, the parameter updates for $\mathcal{A}_{i}$ and $\mathcal{M}_{i}$ are asynchronous. This is due to synchronous updates would result in agents’ private understanding (i.e., $\mathcal{A}_{i}$’s parameters) leaking to their neighborhood before communication through end-to-end training. Thus, the gradient used to update $\mathcal{A}_{i}$ is
$\nabla_{\mathcal{A}_{i}}^{\mathrm{update}}=\nabla_{\mathcal{A}_{i}}\left(J_{\mathcal{A}}(\mathbf{z}^{\mathcal{A}}_{i})\right).$ (4)
To obtain $\nabla_{\mathcal{A}_{i}}^{\mathrm{update}}$, we perform iterative training while freezing the parameters of $\mathcal{M}_{i}$.
III-E2 Decision Rule of $\mathcal{M}_{i}$
During training, parameters of $\mathcal{A}_{i}$ are frozen before updating the parameters of $\mathcal{M}_{i}$, i.e.,
$\nabla_{\mathcal{M}_{i}}^{\mathrm{update}}=\nabla_{\mathcal{M}_{i}}\left(J_{\mathcal{A}}(\mathbf{z}^{\mathcal{A}}_{i})\right).$ (5)
@@ III-E2 Decision Rule of $\mathcal{M}_{i}$
$\mathcal{A}_{i}$ and $\mathcal{M}_{i}$ share the same forward propagation process. Upon explicit request for specific tasks, i.e., when the training of the corresponding model parameters is initiated, $\mathcal{A}_{i}$ and $\mathcal{M}_{i}$ independently perform back propagation based on $J_{\mathcal{A}}(z^{\mathcal{A}}_{i})$, thus maintaining the non-communicability of $\mathcal{A}_{i}$. This concept is illustrated in Fig. 3.
Figure 3: Independent training flow of $\mathcal{A}_{i}$ and $\mathcal{M}_{i}$ with the same forward propagation.
III-E3 Decision Rule of $\mathcal{D}_{i}$
Through the I/O interface, agents obtain their neighbors’ features and their corresponding degrees. Based on this information, the aim of $\mathcal{D}_{i}$ is to learn the relationship between features and degrees from the agent’s limited perspective. In the training phase, the true degrees of neighbors are provided as one-hot encoding to supervise the parameter updates of $\mathcal{D}_{i}$. This implies that the features can be mapped to discrete probability distributions of varying degrees by $\mathcal{D}_{i}$. Furthermore, the agent will allocate weights to the loss corresponding to different neighbors based on the attention vector $W_{i}$, in order to reduce the weight of the loss provided by neighbors with lower attention.
@@ III-E3 Decision Rule of $\mathcal{D}_{i}$
Then, the loss function of $\mathcal{D}_{i}$ is defined as a Kullback-Leibler divergence weighted by neighbor attention, i.e.,
$J_{\mathcal{D}}(\mathcal{X}_{i}^{\mathcal{D}})=-\frac{\mathrm{Sum}\left(\mathcal{X}_{i}^{\mathcal{D}}\log\left(\mathcal{X}_{i}^{\mathcal{D}}\oslash\mathcal{Y}^{\mathcal{D}}_{i}\right)\mathbf{I}\odot\mathbf{w}_{i}\right)}{\mathrm{deg}_{i}+1},$ (8)
where $\mathbf{I}$ is an all-ones vector with $\mathrm{deg}_{\mathrm{max}}$ items that serves to assist in computing the sum of the elements in each row of a matrix.
III-E4 Decision Rule of $\mathcal{N}_{i}$
$v_{i}$ employs $\mathcal{N}_{i}$ to determine whether two nodes have a neighboring relationship based on their features. The training data for $\mathcal{N}_{i}$ is sourced from the sets of features within the scope of $v_{i}$’s view, i.e., $\mathbf{z}_{i}$ and $\{\mathbf{z}_{j}:v_{j}\in\mathrm{N}_{i}\}$. In a graph, the distinguishable decision region exists for the similarity patterns between $v_{i}$ with its first-order and second-order neighbors [47, 48]. Therefore, $\mathcal{N}_{i}$ solely relies on the samples received in the current communication round, to accomplish the binary classification training task between 2 categories: 1) the feature pairs that are labeled as “neighbor” with training samples

### v2

# Graph Agent Network: Empowering Nodes with Inference Capabilities for Adversarial Resilience

@@ Abstract
1 Introduction
Graph neural networks (GNNs) have become state-of-the-art models for node classification tasks by leveraging end-to-end global training paradigms to effectively learn and extract valuable information from graph-structured data (Hamilton et al. 2017). However, this approach has also inadvertently introduced inherent vulnerabilities, making GNNs vulnerable to adversarial edge-perturbing attacks. These vulnerabilities arise from GNNs exposing a global end-to-end training interface, which while allowing for precise classification, also provides adversaries with opportunities to attack GNNs. These attacks poses a significant challenge to the practical deployment of GNNs in real-world applications where security and robustness are of paramount importance, leading to a critical issue in various application areas (Yang et al. 2022), including those where adversarial perturbations undermine public trust (Kreps and Kriner 2020), interfere with human decision making (Walt, Jack, and Christof 2019), and affect human health and livelihoods (Samuel et al. 2019).
@@ 1 Introduction
In response to the edge-perturbing attacks, existing defense mechanisms primarily rely on global-optimization-based defense methods (Sun et al. 2022). Their objective is to enhance the robustness of GNNs through adversarial training (Feng et al. 2019), aiming to increase the tolerant ability of perturbations to defend against potential adversarial perturbations. These approaches may still inherit some inherent vulnerabilities of GNNs, as their model frameworks still follow the global-optimization pattern of GNNs. Representative examples include: (1) RGCN (Zhu et al. 2019) replaces the hidden representations of nodes in each graph convolutional network (GCN) (Kipf and Welling 2017) layer to the Gaussian distributions, to further absorb the effects of adversarial changes. (2) GCN-SVD (Entezari et al. 2020) combines a singular value decomposition (SVD) filter prior to GCN to eliminate adversarial edges in the training set. (3) STABLE (Li et al. 2022) reforms the forward propagation of GCN by adding functions that randomly recover the roughly removed edges. Unfortunately, the computational universality of GNNs has been recently demonstrated (Loukas 2019; Xu et al. 2019), signifying that attributed graphs can be classified into any given label space, even those subject to malicious manipulation. This implies that defense approaches focused on enhancing GNNs’ global robustness (Zhang and Zitnik 2020; Liu et al. 2021; Li et al. 2022; Liu et al. 2024) remain theoretically vulnerable, with potential weaknesses to secondary attacks. In fact, the vulnerability of existing global-optimization-based defenses has been theoretically proven (Liu et al. 2022). Adversaries can reduce the classification accuracy of these defenses once again by launching secondary attacks using these defenses as surrogate models. Experimental evidence demonstrates that even under the protection of state-of-the-art defensive measures (Geisler et al. 2020), secondary attacks targeting these defenses successfully mislead 72.8% of the nodes into making incorrect classifications once again.
@@
To overcome the issues mentioned above, inspired by the natural filtering ability of decentralized intelligence (Saldanha et al. 2022), we propose a decentralized agent network called Graph Agent Networks (GAgN) whose principle is shown in Figure 1. GAgN empowers nodes with autonomous awareness, while limiting their perspectives to their 1-hop neighbors. As a result, nodes no longer rely solely on global-level end-to-end training data. Instead, they progressively gain the perception of the entire network through communication with their neighbors and accomplish self-classification.
Figure 2: The internal structure and communication paradigm of an agent within GAgN.
Specifically, GAgN is designed to enhance the agent-based model (ABM) (Khodabandelu and Park 2021), ensuring improved compatibility with graph-based scenarios. In GAgN, agents are interconnected within the graph topology and engage in cautious communication to mitigate the impact of potential adversarial edges. As time progresses, these agents systematically exchange information to broaden their receptive fields, perceive the global information, and attain decentralized intelligence. An agent primarily includes two major abilities: 1) Storage: states, which are storable features, and actions, which are trainable functions for inference. 2) Communication: agents can receive states and actions from neighboring agents and integrate them to improve their own inference capabilities.
During the communication round, an agent organizes multiple restricted-view mini-datasets based on information received from its immediate neighbors, and updates its states and actions accordingly. After sufficient communications, the agent accumulates enough historical experience to perform specific tasks. These tasks include: 1) computing its own embedding, 2) estimating the possible degree of a given node, and 3) determining the neighboring confidence of two given nodes. The first function enables GAgN to perform node classification, whereas the second and third functions collaboratively contribute to filtering adversarial edges.
@@
Furthermore, we rigorously demonstrate that all these functions can be accomplished by the single-hidden-layer multilayer perceptron (MLPs). In other words, theoretically, within an agent, only three trainable matrices are required in the trainable part, excluding multi-layer networks, bias matrices, and other trainable parameters, to successfully perform the corresponding tasks. This conclusion offers theoretical backing for substantially reducing GAgN’s computational complexity. In practice, by instantiating nodes as lightweight agents, a GAgN can be constructed to carry out node classification tasks with adversarial resilience.
Our main contributions including:
@@
  • We experimentally show that primary functions of GAgN have been effectively executed, which collectively yielding state-of-the-art classification accuracy in perturbed datasets.
2 The Proposed Method
We start with action spaces and discuss the state’s structure and update strategies alongside other components to clarify the entire process. Figure 2 shows the workflow of GAgN.
Preliminaries
@@ 2 The Proposed Method
Consider connected graphs $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with $|\mathcal{V}|$ nodes. The degree of node $v_{i}$ is $\mathrm{deg}_{i}$. The feature vector and one-hot label of $v_{i}$ are $\mathbf{z}_{i}\in\mathbb{R}^{1\times d_{z}}$ and $\mathbf{\ell}_{i}\in\mathbb{R}^{1\times d_{L}}$. The maximum degree is $\mathrm{deg}_{\mathrm{max}}$. The neighborhood $\mathrm{N}_{i}$ of $v_{i}$ includes all $v_{j}$ with $e_{i,j}\in\mathcal{E}$. The edge between $v_{i}$ and $v_{j}$ is $e_{i,j}$. Given a set of vectors, $\mathrm{CONCAT}_{r}(\cdot)$ and $\mathrm{CONCAT}_{c}(\cdot)$ output the row-wise and column-wise concatenated matrices.
2.1 Action Spaces
The action space consists of all potential actions a node can take in a reasoning task (Peng and Van De Panne 2017). In GAgN, it is designed as a set of perception functions with the same parameter structure but different specific parameters. The optimal parameters for agents in each communication round can be found within this action space. We use node $v_{i}$ as an example to illustrate these functions.
Individual Function
Parameters of individual functions are specific to each agent, reflecting the understanding gained through interactions with neighbors, without direct agent interactions. This includes:
@@ 2.1 Action Spaces
$\mathcal{A}_{i}(\mathrm{N}_{i})={\mathbf{w}_{i}\mathrm{Concat}_{r}(\{\mathbf{z}_{i}\}\cup\{\mathbf{z}_{j}:v_{j}\in\mathrm{N}_{i}\})}/{\mathrm{deg}_{i}+1}.$ (1)
$\mathbf{w}_{i}$ is initialized with minimal values and updated iteratively.
Communicable Functions
@@
2.2 I/O Interface
In a single communication round, node $v_{i}$ receives the following messages from its neighbors $\mathrm{N}_{i}$. 1) Communicable functions of all nodes in $\mathrm{N}_{i}$, i.e., sets of functions $\mathcal{M}_{i}^{\mathrm{rec}}=\{\mathcal{M}_{j}:v_{j}\in\mathrm{N}_{i}\}$, $\mathcal{D}_{i}^{\mathrm{rec}}=\{\mathcal{D}_{j}:v_{j}\in\mathrm{N}_{i}\}$, and $\mathcal{N}_{i}^{\mathrm{rec}}=\{\mathcal{N}_{j}:v_{j}\in\mathrm{N}_{i}\}$. 2) Features from the 1-hop neighbors $Z^{\mathrm{I}}_{i}=\{\mathbf{z}_{j}:v_{j}\in\mathrm{N}_{i}\}$. 3) Sampled features from the 2-hop neighbors $Z^{\mathrm{II}}_{i}=\bigcup_{v_{j}\in\mathrm{N}_{i}}\mathrm{S}(v_{j})-Z^{\mathrm{I}}_{i}$, where $\mathrm{S}(\cdot)$ denotes the sample function.
Output
@@ 2.2 I/O Interface
where $\mathrm{rank}(\cdot)$ denotes the rank, a smaller rank indicating a larger value. 3) Communicable functions $\mathcal{M}_{i}$, $\mathcal{D}_{i}$, and $\mathcal{N}_{i}$.
2.3 Decision Rules
@@
The decision rules define how to choose actions in different states (Farmer and Foley 2009). In GAgN, they guide the training of function parameters in the action space, based on $v_{i}$’s states after receiving neighbors’ messages. These functions adjust parameters via the loss function using stochastic gradient descent (SGD).
Figure 3: Independent training flow of $\mathcal{A}_{i}$ and $\mathcal{M}_{i}$.
Aggregator
The loss function $J_{\mathcal{A}}(\cdot)$ of $\mathcal{A}_{i}$ assigns different weights to $v_{i}$’s neighbors to achieve a more accurate embedding closer to its label. For the aggregated feature $\mathbf{z}^{\mathcal{A}}_{i}=\mathcal{A}_{i}(\mathrm{N}_{i})$, the cross-entropy loss is: $J_{\mathcal{A}}(z^{\mathcal{A}}_{i})=-\log\mathcal{M}_{i}(\mathbf{z}^{\mathcal{A}}_{i})\mathbf{\ell}_{i}^{\top}.$ Parameters of $\mathcal{A}_{i}$ are updated via SGD-based backpropagation. Notably, while $\mathcal{M}_{i}$ participates in forward propagation, updates for $\mathcal{A}_{i}$ and $\mathcal{M}_{i}$ occur asynchronously to prevent the leakage of agents’ private understandings (i.e., $\mathcal{A}_{i}$’s parameters) to their neighbors before secure communication through end-to-end training. Consequently, the gradient for updating $\mathcal{A}_{i}$ is $\nabla_{\mathcal{A}_{i}}^{\mathrm{update}}=\nabla_{\mathcal{A}_{i}}\left(J_{\mathcal{A}}(\mathbf{z}^{\mathcal{A}}_{i})\right).$ To obtain $\nabla_{\mathcal{A}_{i}}^{\mathrm{update}}$, we perform iterative training while freezing the parameters of $\mathcal{M}_{i}$.
Embedding Func.
@@
During training, parameters of $\mathcal{A}_{i}$ are frozen before updating the parameters of $\mathcal{M}_{i}$, i.e., $\nabla_{\mathcal{M}_{i}}^{\mathrm{update}}=\nabla_{\mathcal{M}_{i}}\left(J_{\mathcal{A}}(\mathbf{z}^{\mathcal{A}}_{i})\right).$ $\mathcal{A}_{i}$ and $\mathcal{M}_{i}$ share the same forward propagation process. Upon explicit request for specific tasks, i.e., when the training of the corresponding model parameters is initiated, $\mathcal{A}_{i}$ and $\mathcal{M}_{i}$ independently perform back propagation based on $J_{\mathcal{A}}(z^{\mathcal{A}}_{i})$, thus maintaining the non-communicability of $\mathcal{A}_{i}$. This is illustrated in Figure 3.
Inference Func.
Agents use the I/O interface to access neighboring features and degrees. $\mathcal{D}_{i}$ learns to associate these features with degrees. Neighbors’ true degrees, provided as one-hot vectors, guide $\mathcal{D}_{i}$’s updates, mapping features to discrete degree distributions. The attention vector $W_{i}$ adjusts the loss weight for each neighbor, reducing the impact of those with lower attention weights.
@@
Then, the loss function of $\mathcal{D}_{i}$ is defined as a Kullback-Leibler divergence (denoted as $\mathrm{KL}(\cdot)$) weighted by neighbor attention, i.e., $J_{\mathcal{D}}(\mathcal{X}_{i}^{\mathcal{D}})=-\frac{\mathrm{KL}(\mathcal{X}_{i}^{\mathcal{D}}\|\mathcal{Y}^{\mathcal{D}}_{i})\cdot\sum_{j}\mathbf{I}_{j}\mathbf{w}_{i,j}}{\mathrm{deg}_{i}+1},$ where $\mathbf{I}$ is an all-ones vector with $\mathrm{deg}_{\mathrm{max}}$ elements, used to compute the sum of the elements in each row of a matrix.
Neighboring Confidence Func.
$v_{i}$ uses $\mathcal{N}_{i}$ to identify neighboring relationships based on node features. The training data for $\mathcal{N}_{i}$ includes features within $v_{i}$’s view: $\mathbf{z}_{i}$ and $\{\mathbf{z}_{j}:v_{j}\in\mathrm{N}_{i}\}$. In a graph, distinguishable decision regions exist for similarity patterns between $v_{i}$ and its first- and second-order neighbors (Perozzi et al. 2014). Thus, $\mathcal{N}_{i}$ uses samples from the current communication round to train on two categories: 1) feature pairs labeled as ”neighbor”